Question 8.

The Rick Sanchez Corp recently reported $4,678,000 of sales, $3,740,500 of operating costs other than depreciation, and $450,2200 of depreciation. The company had $533,250 of outstanding bonds that carry 7.75% interest rate and its federal plus state income tax rate was 30%. In order to sustain its operations and thus generate future sales and cash flows, the firm was required to spend $1,085,000 to buy new fixed assets and to invest $22,400 in net operating working capital. What was the firm’s free cash flow?

*Please explain step by step, no Excell Spreadsheet*
Answer #1

Interest =7.75%*533250 =41326.88
Tax =(Sales-Operating Expenses-Depreciation-Interest)*Tax Rate =(4678000-3740500-450220-41326.88)*30% =133785.94
Operating cash flow =Sales-Operating Expenses-Taxes =4678000-3740500-133785.94 =803714.06
Free cash Flow=Operating cash flow-New Fixed assets-Net operating working capital =803714.06-1085000-22400 =-303685.94